# Intrinsic Limits of Charge Carrier Mobilities in Layered Halide Perovskites

### v1

@@ Abstract
Layered halide perovskites emerged as a promising class of materials for optoelectronic applications due to their highly tunable physicochemical properties. These materials can be synthesized with different ordering along the stacking axis and number of inorganic layers n, which allows tuning of their optoelectronic properties via both dielectric and quantum confinement effects.[1, 2, 3] Their electronic properties can also be fine-tuned by engeneering the halogen and metal-site composition.[4] Different from many conventional lead-based halide perovskites, the layered perovskites can be highly stable in ambient conditions due to the possibility of introducing hydrophobic spacer cations, allowing significant improvements in their stability while preserving their photo-active properties.[5, 6, 7, 8]
@@
The relaxed GS structures of $n=1$, $n=2$ and three-dimensional models exhibit space-groups Pnma (Number 62), P4/mbm (Number 127), and P4bm (Number 100), respectively. The average Pb-Br bond lengths are nearly identical in all systems ($3.04$, $3.04$, and $3.03$ for three-dimensional, $n=1$, and $n=2$, respectively), while the average in-plane tilting (Angle $\Phi$ in Figure 1) is $153.5$° for CsPbBr3, and $147.5$°, $155.9$° for $n=1$, $n=2$ GS structures, respectively. The structural properties of the three-dimensional compound are consistent with previously reported experimental and theoretical data for CsPbBr3 [23, 24]. In the same manner, the structural parameters obtained for the layered $n=1$ system are in excellent agreement with the experimental data available in the literature. The in-plane tilting angle deviation are of $-2.2\%$, $0.06\%$, $-0.47\%$ and $-0.61\%$ in relation to (PEA)2PbBr4, (MPenDA)PbBr4, (ODA)PbBr4 and (BDA)PbBr4, respectively.[26, 27] The Pb-Br bond length deviation is of $0.98\%$ in relation to (PEA)2PbBr4.[28] The in-plane tilting angles are also consistent with the work of Stoumpos et al., where $n=1$ is shown to exhibit larger Pb-Br-Pb tiltings than $n=2$ and three-dimensional MAPbI3, which have similar tilting angles.[29] This analysis validates our model choice as it confirms that we retain the essential characteristics of the octahedra network and its evolution going from three-dimensional to layered systems.
@@
Figure 2: Phononic and Electronic Structures. a Phonon dispersions and b electronic band structure for the $n=1$ and $n=2$ layered halide perovskites (Csn+1PbnBr3n+1), and the three-dimensional CsPbBr3. The fundamental band gaps and effective masses at band edges are provided.
@@
Having obtained electronic and phonon structures, we can move to the calculation of charge carriers’ mobilities. A first approach is to employ the Drude formula, $\mu=e\tau/m^{*}$, which leads to a mobility ($\mu$) proportional to the carrier lifetime and inversely proportional to the charge carrier’s effective masses ($m^{*}$[31, 32]. By assuming similar carrier lifetimes and comparing the in-plane effective masses obtained within the parabolic approximation (Figure 2), it would appear that the mobility is not drastically affected by the change of dimension. Thus, within these approximations the Drude’s formula cannot explain the observed mobility drop of an order of magnitude when going from three-dimensional to layered halide perovskites.
@@
The ab initio calculated in-plane hole mobilities $\mu$ are shown in Figure 3. The mobilities of all materials are decreasing with the increase of temperature, due to the increase of the charge carriers scattering by lattice vibrations, i.e. electron-phonon interactions. This is by construction the sole scattering mechanism considered in our calculations. Moreover, the in-plane mobility significantly increases when going from the layered to the three-dimensional CsPbBr3 structure, which reveals that dimensionality plays a critical role in the transport properties of perovskites. A similar phenomenon was observed by Li et al. when studying two-dimensional (2D) and bulk InSe [38] and also by Cheng et al. for a variety of 2D semiconductors [39]. The charge carrier mobility increases from 100 cm2/Vs to 1000 cm2/Vs when comparing 2D and bulk InSe. Conversely, carrier mobilities in classic III-V semiconductor quantum wells can reach very high values, thanks to quantum confinement which generate subbands and lift the degeneracy of the valence band.[20, 40] For lead bromide perovskites, at room temperature we predict hole mobilities of $4.3$ cm2/Vs, $15.4$ cm2/Vs and $41.6$ cm2/Vs for the $n=1$, $n=2$ and three-dimensional CsPbBr3 structures, respectively. The obtained mobility for CsPbBr3 is in excellent agreement with the $41$ cm2/Vs previously obtained by Poncé et al. for cubic CsPbBr3.[30] Additionaly, the in-plane electron mobilities for the $n=1$ and $n=2$ layered systems are found to be $8.5$ cm2/Vs and $16.0$ cm2/Vs (Figure S1 of the Support Information file (SI)), respectively, revealing slightly larger mobilities for electrons in agreement with Poncé et al. results for three-dimensional halide perovskites.[30] The increase of charge carrier mobilities as a function of the number of inorganic layers is consistent with previous works by Mitzi et al. [16] and Gélvez-Rueda et al. [3], reporting combined electron-hole mobilities. We also note that for the $n=1$ and $n=2$ systems the mobilites were calculated using the two-dimensional form of the long-range Fröhlich coupling as proposed by W. H. Sio and F. Giustino. [41] In Figure S2 of the SI we show a comparison between the hole mobilities employing the two-dimensional and three-dimensional form of the long-range coupling. This highlights the importance of properly treating the Fröhlich coupling for the two-dimensional compounds, due to the divergency of the el-ph matrix elements for small q.
@@
Figure 3: Average in-plane hole mobilities. Calculations performed within the aiBTE approach for the $n=1$, $n=2$, and three-dimensional systems. The experimental mobilities for the three-dimensional structure MAPbI3 were extracted from reference 19 and the layered $n=1$ mobilities from Reference 18. The experimental values correspond to the combined electron and hole mobilities.
@@
Figure 4: Charge carrier scattering rates, electronic density of states and electron-phonon coupling strength. a Scattering rates evaluated within the aiBTE approach. The points in dark colors represent the average scattering rates for each energy step. b Electronic density of states (DOS) near the valence band edge. The three systems’ VBM are set to zero. c Energy-averaged scattering rate as a function of the electronic density of states.
@@
Figure 5: Spectral decomposition of the contribution of phonons with energy $\hbar\omega$ to the scattering rates $\tau^{-1}$. The spectral decompositions (yellow peaks), and vibrational mode representations are show for the a $n=1$ and b $n=2$ layered, and c three-dimensional halide perovskite. The y-axis of the three-dimensional structure plot is multiplied by fifteen for visibility. The phonon density of states is shown in gray. The cumulative integral in THz is shown in dotted blue. The atoms colored in gold represent the halogens that are not connected along the stacking direction. The gray boxes below each system indicate each system’s space group and $\Gamma$-point point groups. The labels R and I refer to Raman-active and infrared-active, respectively.
@@ Methods
All the calculations within the DFT framework are performed using the Quantum ESPRESSO Suite.[48, 37] The layered structures are optimized with a 80 Ry cutoff for the plane-wave kinetic energy with a $4\times 4\times 1$ Brillouin zone sampling while a $4\times 4\times 4$ Brillouin zone sampling was employed for the three-dimensional structure. The threshold on forces and total energy during ionic minimization were converged to $10^{-6}$ and $10^{-11}$, respectively. During the optimizations a 0.1 Kbar threshold is employed for the pressure on the variable cell. A threshold of $10^{-11}$ is employed for the SCF steps. All the calculations for hole transport properties are performed using non-relativistic norm-conserving PBE pseudopotential taken from the pseudo-dojo database (http://www.pseudo-dojo.org) [49], while electron transport properties are evaluated taking spin-orbit interactions into account. The phonon calculations are performed within the DFPT framework as also implemented in Quantum ESPRESSO. A $4\times 4\times 4$ k-grid is employed for the SCF, while the phonons are solved in a $4\times 4\times 1$ q-grid for layered systems and $2\times 2\times 2$ q-grid for the three-dimensional system. A 10-17 threshold for phonon self-consistency is employed.
@@
The solution of the aiBTE was performed with the EPW code.[34, 35, 36, 33, 50] For the calculation of layered systems we employ a $8\times 8\times 1$ and $4\times 4\times 1$ coarse k-grid and q-grid, respectively. These are interpolated in fine grids of $60\times 60\times 1$ for both k-grid and q-grid. The interpolation of electronic band structure is performed using Wannier interpolation. For the interpolation of valence bands we employ Pb s-orbitals only, while conduction bands are interpolated using Pb p-orbitals. All states within 300 meV from the band edges are included in the aiBTE calculation. For the three-dimensional system, we employ a $4\times 4\times 4$ and $2\times 2\times 2$ coarse k-grid and q-grid, respectively. These are interpolated in fine grids of $50\times 50\times 40$ for both k-grid and q-grid. For the Wannier interpolation Pb s-orbitals and p-orbitals are employed. All states within 300 meV from the band edges are included in the aiBTE calculation.
@@ acknowledgments
We gratefully acknowledge the discussions with Dr. Hyungjun Lee and Prof. S. Thébaud. The research leading to these results was supported by the Agence Nationale pour la Recherche through the CPJ program and the SURFIN project (ANR-23-CE09-0001). This work was granted access to the HPC resources of TGCC under the allocation 2022-A0130907682 made by GENCI. B.C, J.L, V.H. and F.G. are supported by the Robert A. Welch Foundation under Award Number F-2139-20230405 (calculations and manuscript preparation) and the Computational Materials Sciences Program funded by the U.S. Department of Energy, Office of Science, Basic Energy Sciences, under Award No. DE-SC0020129 (EPW software development). The authors acknowledge the Texas Advanced Computing Center (TACC) at The University of Texas at Austin for providing HPC resources, including the Lonestar6 and Frontera under LRAC award 2103991, that have contributed to the research results reported within this paper. This research also used resources of the National Energy Research Scientific Computing Center, a DOE Office of Science User Facility supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC02-05CH11231.

### v2

@@ Abstract
Layered halide perovskites emerged as a promising class of materials for optoelectronic applications due to their highly tunable physicochemical properties. These materials can be synthesized with different ordering along the stacking axis and number of inorganic layers n, which allows tuning of their optoelectronic properties via both dielectric and quantum confinement effects.[1, 2, 3] Their electronic properties can also be fine-tuned by engineering the halogen and metal-site composition.[4] Different from many conventional lead-based halide perovskites, the layered perovskites can be highly stable in ambient conditions due to the possibility of introducing hydrophobic spacer cations, allowing significant improvements in their stability while preserving their photo-active properties.[5, 6, 7, 8]
@@
The relaxed GS structures of $n=1$, $n=2$ and three-dimensional models exhibit space-groups P4/mbm (Number 127), P4bm (Number 100) and Pnma (Number 62), respectively. The average Pb-Br bond lengths are nearly identical in all systems ($3.04$, $3.04$, and $3.03$ for three-dimensional, $n=1$, and $n=2$, respectively), while the average in-plane tilting (Angle $\Phi$ in Figure 1) is $153.5$° for CsPbBr3, and $147.5$°, $155.9$° for $n=1$, $n=2$ GS structures, respectively. The structural properties of the three-dimensional compound are consistent with previously reported experimental and theoretical data for CsPbBr3 [23, 24]. In the same manner, the structural parameters obtained for the layered $n=1$ system are in excellent agreement with the experimental data available in the literature. The in-plane tilting angle deviation are of $-2.2\%$, $0.06\%$, $-0.47\%$ and $-0.61\%$ in relation to (PEA)2PbBr4, (MPenDA)PbBr4, (ODA)PbBr4 and (BDA)PbBr4, respectively.[26, 27] The Pb-Br bond length deviation is of $0.98\%$ in relation to (PEA)2PbBr4.[28] The in-plane tilting angles are also consistent with the work of Stoumpos et al., where $n=1$ is shown to exhibit larger Pb-Br-Pb tiltings than $n=2$ and three-dimensional MAPbI3, which have similar tilting angles.[29] This analysis validates our model choice as it confirms that we retain the essential characteristics of the octahedra network and its evolution going from three-dimensional to layered systems.
@@
Figure 2: Phononic and Electronic Structures. (a) Phonon dispersions and (b) electronic band structure for the $n=1$ and $n=2$ layered halide perovskites (Csn+1PbnBr3n+1), and the three-dimensional CsPbBr3. The fundamental band gaps and effective masses at band edges are provided.
@@
Having obtained electronic and phonon structures, we can move to the calculation of charge carriers’ mobilities. To provide insights on how the change in the effective mass between these systems affects the charge carrier mobilities, we first employ the Drude’s formula, $\mu=e\tau/m^{*}$, which leads to a mobility ($\mu$) proportional to the carrier lifetime and inversely proportional to the charge carrier’s effective masses ($m^{*}$). [31, 32] By assuming similar carrier lifetimes and comparing the in-plane effective masses obtained within the parabolic approximation (Figure 2), it would appear that the mobilities are not drastically affected by the changes in the effective masses originated from the dimensionality reduction. Thus, within these approximations, we conclude that by accounting only changes in the effective masses, Drude’s formula cannot explain the observed mobility drop of an order of magnitude when going from three-dimensional to layered halide perovskites.
@@
The ab initio calculated in-plane hole mobilities $\mu$ are shown in Figure 3. The mobilities of all materials are decreasing with the increase of temperature, due to the increase of the charge carriers scattering by lattice vibrations, i.e. electron-phonon interactions. This is by construction the sole scattering mechanism considered in our calculations. Moreover, the in-plane mobility significantly increases when going from the layered to the three-dimensional CsPbBr3 structure, which reveals that dimensionality plays a critical role in the transport properties of perovskites. A similar phenomenon was observed by Li et al. when studying two-dimensional (2D) and bulk InSe [38] and also by Cheng et al. for a variety of 2D semiconductors [39]. The charge carrier mobility increases from 100 cm2/Vs to 1000 cm2/Vs when comparing 2D and bulk InSe. Conversely, carrier mobilities in classic III-V semiconductor quantum wells can reach very high values, thanks to quantum confinement which generate subbands and lift the degeneracy of the valence band.[20, 40] For lead bromide perovskites, at room temperature we predict hole mobilities of $4.3$ cm2/Vs, $15.4$ cm2/Vs and $41.6$ cm2/Vs for the $n=1$, $n=2$ and three-dimensional CsPbBr3 structures, respectively. The obtained mobility for CsPbBr3 is in excellent agreement with the $41$ cm2/Vs previously obtained by Poncé et al. for cubic CsPbBr3 also within the same fully ab initio approach.[30] Other sets of calculations for CsPbX3 (X = Cl, Br, I) employing different methodologies can also be found in the literature, such as Feynman polaron models [41], Bloch–Boltzmann theory[42] and numerical integration of the BTE using LO electron-phonon matrix elements within the Fröhlich model [43]. Additionally, the in-plane electron mobilities for the $n=1$ and $n=2$ layered systems are found to be $8.5$ cm2/Vs and $16.0$ cm2/Vs (Figure S1 of the Support Information file (SI) [44]), respectively, revealing slightly larger mobilities for electrons in agreement with Poncé et al. results for three-dimensional halide perovskites.[30] It has been shown that the well-known DFT underestimation of the band-gap values, can lead to a slight overestimation of mobilities of up to 15%.[45] The increase of charge carrier mobilities as a function of the number of inorganic layers is consistent with previous works by Mitzi et al. [16] and Gélvez-Rueda et al. [3], reporting combined electron-hole mobilities. We also note that for the $n=1$ and $n=2$ systems the mobilites were calculated using the two-dimensional form of the long-range Fröhlich coupling as proposed by W. H. Sio and F. Giustino. [46] In Figure S2 of the SI we show a comparison between the hole mobilities employing the two-dimensional and three-dimensional form of the long-range coupling. [44] This highlights the importance of properly treating the Fröhlich coupling for the two-dimensional compounds, due to the divergency of the el-ph matrix elements for small q.
@@
Figure 3: Average in-plane hole mobilities. The calculations performed within the aiBTE approach for the $n=1$, $n=2$, and three-dimensional systems are shown as solid lines. The experimental mobilities for the three-dimensional structures CsPbBr3 and MAPbI3 were extracted from references 47 and 19, respectively. The mobility for (PEA)2PbI4 was extracted from Reference 18. The experimental values correspond to the combined electron and hole mobilities.
@@
Figure 4: Charge carrier scattering rates, electronic density of states and electron-phonon coupling strength. (a) Scattering rates evaluated within the aiBTE approach. The points in dark colors represent the average scattering rates for each energy step. (b) Electronic density of states (DOS) near the valence band edge. The three systems’ VBM are set to zero. (c) Energy-averaged scattering rate as a function of the electronic density of states.
@@
Figure 5: Spectral decomposition of the contribution of phonons with energy $\hbar\omega$ to the scattering rates $\tau^{-1}$. The spectral decompositions (yellow peaks), and vibrational mode representations are show for the $n=1$ and $n=2$ layered, and three-dimensional halide perovskite. The y-axis of the three-dimensional structure plot is multiplied by fifteen for visibility. The phonon density of states is shown in gray. The cumulative integral in THz is shown in dotted blue. The atoms colored in gold represent the halogens that are not connected along the stacking direction. The gray boxes below each system indicate each system’s space group and $\Gamma$-point point groups. The labels R and I refer to Raman-active and infrared-active, respectively.
@@ Methods
All the calculations within the DFT framework are performed using the Quantum ESPRESSO Suite.[54, 37] The layered structures are optimized with a 80 Ry cutoff for the plane-wave kinetic energy with a $4\times 4\times 1$ Brillouin zone sampling while a $4\times 4\times 4$ Brillouin zone sampling was employed for the three-dimensional structure. The threshold on forces and total energy during ionic minimization were converged to $10^{-6}$ Ry/bohr and $10^{-11}$ Ry, respectively. During the optimizations, a 0.1 Kbar threshold is employed for the pressure on the variable cell. A threshold of $10^{-11}$ Ry is employed for the SCF steps. All the calculations for hole transport properties are performed using non-relativistic norm-conserving PBE pseudopotential taken from the pseudo-dojo database (http://www.pseudo-dojo.org) [55], while electron transport properties are evaluated taking spin-orbit interactions into account. The self-interaction error due to the DFT-PBE exchange-correlation functional has marginal effects on the computation of the electron-phonon matrix elements, as the charge carriers are delocalized.[25] Furthermore, although in perovskites the inclusion of spin-orbit coupling is fundamental to describe the complete valence band manifold, in the vicinity of the band edges there are no significant changes in energy and band dispersion, as shown in Figure S3. [44] The phonon calculations are performed within the DFPT framework as also implemented in Quantum ESPRESSO. A $4\times 4\times 4$ k-grid is employed for the SCF, while the phonons are solved in a $4\times 4\times 1$ q-grid for layered systems and $2\times 2\times 2$ q-grid for the three-dimensional system. A 10-17 Ry threshold for phonon self-consistency is employed.
@@
The solution of the aiBTE was performed with the EPW code.[34, 35, 36, 33, 56] For the calculation of layered systems we employ a $8\times 8\times 1$ and $4\times 4\times 1$ coarse k-grid and q-grid, respectively. These are interpolated in fine grids of $60\times 60\times 1$ for both k-grid and q-grid. The interpolation of the electronic band structure is performed using Wannier interpolation. For the Wannier interpolation of the valence and conduction band, we start with only Pb orbitals as an initial set of projectors. The resulting orbitals lead to high-quality interpolation in the narrow energy range relevant to transport. The Wannier interpolated band structure and DFT, along with the corresponding MLWFs can be found in Figure S4 of the SI. [44] We note that a more complete set would be required to interpolate the complete valence and conduction band manifold. All states within 300 meV from the band edges are included in the aiBTE calculation. For the three-dimensional system, we employ a $4\times 4\times 4$ and $2\times 2\times 2$ coarse k-grid and q-grid, respectively. These are interpolated in fine grids of $50\times 50\times 40$ for both k-grid and q-grid. For the Wannier interpolation Pb s-orbitals and p-orbitals are employed. All states within 300 meV from the band edges are included in the aiBTE calculation.
@@ acknowledgments
We gratefully acknowledge the discussions with Prof. V. Podzorov and Dr. V. Bruevich regarding intrinsic mobility measurements, and the fruitfull discussions with Dr. Hyungjun Lee and Prof. S. Thébaud. The research leading to these results was supported by the Agence Nationale pour la Recherche through the CPJ program and the SURFIN project (ANR-23-CE09-0001). This work was granted access to the HPC resources of TGCC under the allocation 2022-A0130907682 made by GENCI. B.C, J.L, V.H. and F.G. are supported by the Robert A. Welch Foundation under Award Number F-2139-20230405 (calculations and manuscript preparation) and the Computational Materials Sciences Program funded by the U.S. Department of Energy, Office of Science, Basic Energy Sciences, under Award No. DE-SC0020129 (EPW software development). The authors acknowledge the Texas Advanced Computing Center (TACC) at The University of Texas at Austin for providing HPC resources, including the Lonestar6 and Frontera under LRAC award 2103991, that have contributed to the research results reported within this paper. This research also used resources of the National Energy Research Scientific Computing Center, a DOE Office of Science User Facility supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC02-05CH11231.